The probability that an event happens in one trial of an experiment, is 0.4. Three independent trials of the experiments are performed. The probability that the event happens atleast once, is
- A.0.936
- B.0.784
- C.0.904
- D.None of these
Show answer and solution
Correct option: B
Solution
P(A doesn't happen in one trial) = 1 - 0.4 = 0.6
P(A doesn't happen in all three trials) = (0.6)³ = 0.216
P(A happens at least once) = 1 - 0.216 = 0.784
P(A doesn't happen in all three trials) = (0.6)³ = 0.216
P(A happens at least once) = 1 - 0.216 = 0.784
[1980]